OpenStudy (anonymous):

find the limit: lim x>4 sqrt x-2/x-4

ganeshie8 (ganeshie8):

hint : \(\large x - 4 = (\sqrt{x})^2 - 2^2 = \left(\sqrt{x}+2\right)\left(\sqrt{x}-2\right)\)